Solve Laplace's Equation nabla^2 V(z)=0.

1
Expert's answer
2022-01-31T07:34:33-0500

To solve Laplace's equation


"\nabla^2 V(z)=0,"

we need to know how the function "V(z)" looks. However, the solution can be written as


"\nabla^2V(z)=\dfrac{\partial^2 V(z)}{\partial z^2}=0,\\\space\\
V(z)=mz+c,"

where m and c are constants defined by boundary conditions.